"Find sin theta/2 given that sin theta=-15/17,
where pi is less than theta is less than 3pi/2."

Answer by stanbon(75887)

"Find sin theta/2 given that sin theta=-15/17,
where pi is less than theta is less than 3pi/2."
----------------------------------------------
Form:: sin(x/2) = sqrt[(1-cos(x))/2
-----
Since sin = y/r, y = -15 ; r = 17
-----
In QIII x is negative.
-----
x = -sqrt[17^2-15^2] = -8
---
Then cos(t) = -8/17
----------------------------
Ans: sin(t/2) = sqrt[(1-(-8/17))/2] = sqrt[25/34] = 5/sqrt(34)
=================================================================
Cheers,
Stan H.
=====================